OpenStudy (anonymous):

will gave a medal What is the equation of the line that passes through the points (–2, 3) and (2, 7)? A. x – y = –1 B. x – y = –2 C. x – y = –5 D. x – y = –6

OpenStudy (triciaal):

you have 2 points find the slope choose one point you have y,m and x find b now write the general equation in the form y =mx + b

OpenStudy (triciaal):

sorry in this problem they want x - y =b

OpenStudy (anonymous):

so -2-3=-1?

OpenStudy (triciaal):

how do you find the slope

OpenStudy (triciaal):

change in y/change in x 7-3/2--2 =4/4 =1 using the point (2 7) 7 = (1)(2) + b b=5 y = x + 5 rearrange to get x - y = ?

OpenStudy (triciaal):

are you here